OpenStudy (anonymous):

A large square carpet is initially x inches long and x inches wide. Randy decides to trim 2 inches off each edge of the carpet because the edges are worn. Later he goes back and trims 6 more inches off each edge. (a) Express the final area (A) of the carpet in square inches (after both trim jobs) as a polynomial function of x. A(x)= (b) How many square inches of material were removed in the first trim job? Express the area removed in the first trim job (T1) as a function of x. T1(x)= (c) How many square inches of material were removed in the second trim job? Express the area removed in t

OpenStudy (anonymous):

im a fan brandon

OpenStudy (anonymous):

a fan of you. if u answer this

OpenStudy (anonymous):

Well I guess since you put me on the spot :P

OpenStudy (anonymous):

area=x^2 area after first trimming=(x-2)^2 area after second trimming=(x-8)^2 a. A(x)=x^2-16x+64 b. T1(x)=x^2-(x-2)^2 c. T2(x)=x^2-(x-8)^2